A radiator contains 12 quarts of fluid, 35% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 60% antifreeze? (Round your answer to two decimal places.)
Answer by richwmiller(17219)

0.35*12-0.35x+1x=0.6(12)
4.2+0.65x=7.2
0.65x=7.2-4.2
0.65x=3
x=4.61538462
check
0.35*7.38461538+1*4.61538462=0.6(12)
2.58461538+4.61538462=7.2
7.2=7.2
ok
codemixrep